OpenStudy (anonymous):

The length of a rectangle is six times its width. If the perimeter of the rectangle is 98 yd , find its area.

OpenStudy (anonymous):

Let's call the rectangle u The sides of the rectangle are a and b: Then is \[u = 2 \cdot (a+b)\] And \[a=6 b\] \[u = 2 \cdot (6b+b) = 14 \cdot b\] SO the width is 7 and the lengt is 42

OpenStudy (anonymous):

And the area is 7x42?

OpenStudy (anonymous):

Yes of course, what else?

OpenStudy (anonymous):

OoohThanks :)